OpenStudy (anonymous):

Helpppp plzzzzz Part 1: Simplify f(x) = the quantity 2 x-squared plus 3x plus 1 all over 2x plus 1. (4 points) Part 2: Using complete sentences, explain why f(1) = 2, f(0) = 1, and f(-1) = 0, yet f(-one¨Chalf) is undefined. Be sure to show your work. (4 points)

OpenStudy (campbell_st):

well you need to factor the quadratic in the numerator... there is a common factor \[\frac{2x^2 + 3x + 1}{2x + 1}\]

OpenStudy (anonymous):

Ah ok not sure how to do that but ok lol thanks

OpenStudy (campbell_st):

well the numerator becomes \[\frac{(2x + 1)(x + 1)}{(2x + 1)}\]

OpenStudy (campbell_st):

for part 2 substitute x = 1 into the expression and get an answer then x = 0 get the answer x = -1 get the answer for x = -1/2 show that when you substitute it into the denominator you get zero... any fraction is undefined when a zero denominator occurs... this is the restriction... hope it helps

OpenStudy (anonymous):

ah cool thanks Helps a lot